OpenStudy (anonymous):

CAN SOMEONE HELP ME PLEASE The formula for the perimeter of a rectangle is P = 2l + 2w. Find w if P = 68 and l = 10. 25 15 28 24

OpenStudy (anonymous):

68=2(l+w) then 34=l+w 34-10=W W=24 unit